OpenStudy (anonymous):

Write an equation for the line in point/slope form and slope/intercept form that has the given condition. Slope = 1/2 and passes through the origin. I kinda forgot how to do this, can someone jog my memory please?

OpenStudy (jhannybean):

you know your equation will end up in slope intercept form, \[\large y=mx+b\] after deriving it from point-slope form, yes?

OpenStudy (jdoe0001):

http://www.teachhandson.com/calculationb.gif

OpenStudy (jhannybean):

We start with point-slope form, with your points (0,0) \[(x_{1},y_{1}) = (0,0)\] and slope \(m=\frac12\). \[y-y_{1}=m(x-x_{1})\\y-(0)=\frac12(x-(0))\\y=\frac12x\]

OpenStudy (anonymous):

@Jhannybean y=1/2x is the point-slope form?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

y=1/2x would be the point-slope form y=mx+b would be the slope intercept form, now how do I plug in to make it fit the problem ?